OpenStudy (anonymous):

A sphere is inscribed in a right circular cylinder . What is the ratio if the volume of the sphere to the volume of the cylinder?

OpenStudy (anonymous):

Do you know the formulas for the volume of spheres and cylinders?

OpenStudy (anonymous):

Yes, V =πr^2h Cylinder. V = 4^3 π r Sphere

OpenStudy (anonymous):

 Since the sphere is inscribed within the cylinder, the sphere's diameter is equal to both the cylinder's height, and it's diamter

OpenStudy (anonymous):

okay… How do we find the ratio through the formals. I can do it with numbers. but without I get really confused.

OpenStudy (anonymous):

Okay, so you have these two formulas for volume: \[V_s = 4/3*\pi*(r_s)^3 \] \[V_C = \pi*(r_C)^2*h_C\] We know that \[D_s = D_C\] and \[D_s = h_C\] Also, the diameter is equal to the radius times two, so \[r_C = D_C/2 \] and \[r_s = D_s/2\] Now that you can relate the dimensions of the cylinder to the sphere, you can take the ratio of the sphere's volume to the cylinder's volume: \[ratio = V_s/V_C \]

OpenStudy (anonymous):

As a hint, try to find rC and hC in terms of rs. That way, when you divide to find the ratio, the rs terms will all cancel out and you'll be left with a constant, which will be your ratio

OpenStudy (anonymous):

I think I got it now! Thanks for your help! :)